#Python I have a data frame called data_tweets, there is one column calls timestamp, the dataset looks like: 2016-06-25...

#Python

I have a data frame called data_tweets, there is one column calls timestamp, the dataset looks like: 2016-06-25 23:59:52

And there is another data frame called data_BTUSD, the time column looks like:27/5/19 0:00

How can I make these two datafram's time in the same format and draw the time line for them?

Thanks!
Answer #1

Now first you have to extract the column timestamp from data frame data_tweets in a vector time1 and then you have to extract the column time from data frame data_BTUSD in a vector time2 now use the below program to convert time1 values to time2 and then draw a timeline for them:

PYTHON CODE:

def compare(t1,t2):
    if t1[1]=='/':
        day1=int(t1[0])
        if t1[3]=='/':
            month1=int(t1[2])
            year1=int(t1[4:6])
        else:
            month1=int(t1[2:4])
            year1=int(t1[5:7])
    else:
        day1=int(t1[0:2])
        if t1[4]=='/':
            month1=int(t1[3])
            year1=int(t1[5:7])
        else:
            month1=int(t1[3:5])
            year1=int(t1[6:8])
    if t2[1]=='/':
        day2=int(t2[0])
        if t2[3]=='/':
            month2=int(t2[2])
            year2=int(t2[4:6])
        else:
            month2=int(t2[2:4])
            year2=int(t2[5:7])
    else:
        day2=int(t2[0:2])
        if t2[4]=='/':
            month2=int(t2[3])
            year2=int(t2[5:7])
        else:
            month2=int(t2[3:5])
            year2=int(t2[6:8])
    ind1=0
    ind2=0
    for i in range(0,len(t1)):
        if t1[i]==' ':
            ind1=i
            break
    for i in range(0,len(t2)):
        if t2[i]==' ':
            ind2=i
            break
    if t1[ind1+2]==':':
        hour1=int(t1[ind1+1])
        minute1=int(t1[ind1+3:ind1+5])
    else:
        hour1=int(t1[ind1+1:ind1+3])
        minute1=int(t1[ind1+4:ind1+6])
    if t2[ind2+2]==':':
        hour2=int(t2[ind2+1])
        minute2=int(t2[ind2+3:ind2+5])
    else:
        hour2=int(t2[ind2+1:ind2+3])
        minute2=int(t2[ind2+4:ind2+6])
    if year1>year2:
        return 1
    elif year1<year2:
        return 0
    else:
        if month1>month2:
            return 1
        elif month1<month2:
            return 0
        else:
            if day1>day2:
                return 1
            elif day1<day2:
                return 0
            else:
                if hour1>hour2:
                    return 1
                elif hour1<hour2:
                    return 0
                else:
                    if minute1>minute2:
                        return 1
                    elif minute1<minute2:
                        return 0
                    else:
                        return 1
#5 values in time1
time1=['2016-06-25 23:59:52','2018-07-30 00;46:12','2019-05-27 06:00:19','2019-05-27 15:11:26','2020-11-21 01:00:11']
#5 values in time2
time2=['27/5/19 0:00','27/5/19 13:13','28/5/19 4:50','30/7/18 0:45','30/7/19 7:34']
for i in time1:
    t=""
    t+=i[8:10]
    t+='/'
    if i[5]=='0':
        t+=i[6]
    else:
        t+=i[5:7]
    t+='/'
    t+=i[2:4]
    t+=' '
    if i[11]=='0':
        t+=i[12]
    else:
        t+=i[11:13]
    t+=":"
    t+=i[14:16]
    time2.append(t)
print("TIME1 values appended in TIME2")
print(time2)
for i in range(0,len(time2)):
    for j in range(i+1,len(time2)):
        if compare(time2[i],time2[j]):
            time2[i],time2[j]=time2[j],time2[i]
print("Drawing Timeline: ")
for i in time2:
    print(i)

Output:

Run test9 Aahi ah Shr ycharmProjects Scripts\venv\Scripts\python.exe C:/Users/Ashish Shrivastav/PycharmProjects/Scripts/tes